a) A company makes tablets of which 85% of tablets would pass a quality control test by an inspector. If an inspector was to pick six tablets at random from a batch and at least four must pass in order for the batch to be approved, what is the probability that the batch passes inspection?
Answer by vashistrs76(1)

The probability that the batch passes inspection: 0.9526
Solution:
Number of tablets that pass the inspection, X = 4, 5, 6
n = 6
p = 0.85
q = 1-85 = 0.15
P (x ≥ 4) = P (x =4) + P(x = 5) + P(x = 6)
P (x = r) = (nCr) (p)^r q^(n-r)
P (x = 4) = (6C4) (0.85)^4 (0.15)^(6-4)
P (x = 4) = 0.1762
P (x = 5) = (6C5) (0.85)^5 (0.15)^(6-5)
P (x = 5) = 0.3993

P (x = 6) = (6C6) (0.85)^6 (0.15)^(6-6)
P (x = 6) = 0.3771

P (x ≥ 4) = 0.1762 + 0.3993 + 0.3771
P (x ≥ 4) = 0.9526
Hence, there is a 0.9526 probability that the batch passes inspection.